# Computable Thermodynamic Formalism

### v1

@@ Abstract.
We investigate the theory of thermodynamic formalism from the perspective of computable analysis, with a special focus on the computability of equilibrium states. Specifically, we develop two complementary general approaches to verify the computability of equilibrium states for non-uniformly expanding computable dynamical systems. The first approach applies to dynamical systems whose topological pressure functions admit effective approximations and whose measure-theoretic entropy functions are upper semi-continuous. As a concrete application, we establish the computability of the equilibrium states for Misiurewicz–Thurston rational maps with Hölder continuous potentials. The second approach exploits prescribed Jacobians of equilibrium states through a local analysis and applies to settings where the measure-theoretic entropy functions may lack upper semi-continuity.
@@ 1. Introduction
Thermodynamic formalism emerges as the natural framework for studying chaotic dynamics from a statistical viewpoint. This theory, which draws inspiration from statistical mechanics, was pioneered by Ruelle, Sinai, Bowen, and others in the 1970s ([Do68, Si72, Bo75, Wa82]). Since its birth, thermodynamic formalism has been extensively applied in various classical contexts and has remained at the frontier and core of research in dynamical systems. It focuses on equilibrium states, which are invariant measures that maximize pressure functionals encoding both entropy [^1] The concepts of entropy in dynamical systems have their roots in the early works on the notions of entropy by Boltzmann and Gibbs (statistical mechanics 1875), von Neumann (quantum mechanics 1932), and Shannon (information theory 1948). These notions of entropy are all designed to describe the complexity of their respective systems or objects. In recent years, there have been exciting developments and diverse applications of entropy and complexity theory, see e.g. Braverman’s report at the International Congress of Mathematicians in 2022 [Bra23]. and integrals of potentials. Crucially, equilibrium states characterize the typical behavior of the dynamics and thus possess central mathematical and physical significance. In many settings, equilibrium states describe the weighted distribution of random backward orbits (see e.g. [HT03, Li18]), iterated preimages, and periodic orbits.
@@
In this article, we study thermodynamic formalism from the point of view of computable analysis, with a special focus on the computability of equilibrium states in dynamical systems. Classical hyperbolic systems admit well-developed techniques and are generally regarded as well-understood. By contrast, non-uniformly hyperbolic systems resist conventional approaches. The relaxation of uniform expansion requirements therefore represents a fundamental challenge in dynamical systems. Our central contribution lies in demonstrating that statistical computability can coexist with dynamical complexity, even in non-uniformly hyperbolic regimes.
@@
The first approach utilizes the set of tangent functionals to the topological pressure function at the potential to link the computability of the topological pressure to that of the equilibrium state. This method is suitable for dynamical systems where the measure-theoretic entropy map is upper semi-continuous and the topological pressure function can be effectively approximated. For dynamical systems whose measure-theoretic entropy functions may lack upper semi-continuity, we implement the second approach, which establishes the computability of equilibrium states through local (rather than global) considerations. The novelties of our second approach include the following two aspects: (i) on the thermodynamic formalism side, using general transfer operators to study equilibrium states of full measure away from singular points with prescribed Jacobians, and (ii) on the computable analysis side, the verification of Jacobians away from singular points, and the construction of a recursively compact set of measures to exclude atomic measures supported on singular points.
@@
Based on these works, ergodic theory for expanding Thurston maps has been actively investigated in [BM10, BM17, HP09, Li15, Li17, Li18, LZ25, LSZ25, LS24a, LS24b]. In particular, the third-named author [Li18] developed the thermodynamic formalism and investigated the existence, uniqueness, and other ergodic properties of equilibrium states for expanding Thurston maps333cf. the monograph [Li17].. Notably, expanding Thurston maps exhibit weak expansion; they are neither expansive nor $h$-expansive. Furthermore, those with at least one periodic critical point are not even asymptotically $h$-expansive [Li15]. Recent advances by the third-named and fourth-named authors [LS24b] further demonstrated that the measure-theoretic entropy function is upper semi-continuous if and only if the expanding Thurston map has no periodic critical points. Leveraging these results, we apply our methods to investigate the computability of equilibrium states for expanding Thurston maps, demonstrating the applicability of our two distinct approaches.
@@ 1.1. Main results
The first approach (Theorem 1.1) applies to systems with upper semi-continuous entropy. It shows that certain computability properties of the topological pressures guarantee the computability of the equilibrium states, with a concrete application to Misiurewicz–Thurston rational maps in Theorem 1.2. The second approach, established in Theorem 1.3, applies to systems without upper semi-continuous entropy, where the computability of equilibrium states arises from the computability of their prescribed Jacobians. As an application, we establish the computability of the measures of maximal entropy for computable expanding Thurston maps with computable critical points in Theorem 1.4.
@@
We say that the quadruple $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ is a uniformly computable system if the following properties are satisfied:
@@
In the first approach, we exploit the set of tangent functionals to the topological pressure function $P_{T}$ (see Section 4) at the potential $\phi$ and demonstrate that the computability of the equilibrium state follows from the certain computability properties of the topological pressure. We note that this set of tangent functionals is naturally identified with a subset of the set $\mathcal{P}(X)$ of Borel probability measures (see Remark 4.4). We denote by $\mathcal{E}(T,\phi)$ the set of equilibrium state(s) for a map $T$ and a potential $\phi$. We refer the reader to Subsection 3.1 for a detailed discussion on the computability of measures.
@@
Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system with $X_{n}=X$ for all $n\in\mathbb{N}$, and $\{\phi_{n}\}_{n\in\mathbb{N}}$ be a sequence of uniformly computable functions $\phi_{n}\colon X\rightarrow\mathbb{R}$. Suppose $T_{n}$ has finite topological entropy, and the measure-theoretic entropy map $\nu\mapsto h_{\nu}(T_{n})$ is upper semi-continuous on $\mathcal{M}(X,T_{n})$ for each $n\in\mathbb{N}$. Assume that the following statements are true:
@@
Here the computability properties of real numbers is given in Definitions 3.3 and 3.6, and the computable structure on the space $\mathcal{P}(X)$ of Borel probability measures is specified in Proposition 3.26.
Applying Theorem 1.1, we establish the computability of the equilibrium states for computable Misiurewicz–Thurston rational maps (i.e., postcritically-finite rational maps without periodic critical points which are computable) and computable Hölder continuous potentials. The existence and the uniqueness of the equilibrium states for expanding Thurston maps follow from Theorem 6.3.
@@
The computable structure in $\widehat{\mathbb{C}}$ is given in Proposition 6.5. Moreover, $\mathopen{\lvert}\phi\mathclose{\rvert}_{\alpha,\sigma}$ denotes the Hölder constant of $\phi\in C^{0,\alpha}\bigl(\widehat{\mathbb{C}},\sigma\bigr)$ (see (2.1)).
It is worth noting that Theorem 1.1 can be applied to establish the computability of the equilibrium states for a wide range of dynamical systems with a unique equilibrium state and an upper semi-continuous measure-theoretic entropy map, such as rational maps with Hölder continuous hyperbolic potentials (see e.g. [DU91]). Due to space limitations, we focus on the current examples and postpone further investigations to future works.
For many dynamical systems, the measure-theoretic entropy map may not be upper semi-continuous, or verifying this property may be difficult. To overcome this issue, we establish a second approach by considering the prescribed Jacobians for equilibrium states.
@@
By Theorem 5.6, in some settings (cf. Definition 5.2), the set defined in (1.2) could describe the set of equilibrium states. Moreover, as applications of Theorem 1.3, in Theorems 5.9 and 5.10, we construct two corresponding families of uniformly recursively compact sets $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ satisfying the additional assumptions for two classes of dynamical systems.
Theorem 1.3 extends the methodologies developed in [BBRY11, Theorem A] and [BHLZ25, Theorem 1.1], introducing new techniques to handle both non-uniform expansion and the failure of upper semi-continuity of entropy. Compared to the previous results, our approach significantly broadens applicability by relaxing the computability requirement on prescribed Jacobians to only lower semi-computability.
To demonstrate the applicability of Theorem 1.3 while minimizing unnecessary technicalities, we employ it to prove the computability of the measures of maximal entropy for computable expanding Thurston maps.
@@
We provide examples of such expanding Thurston maps and apply Theorem 1.4 to them at the end of Subsection 6.3. It is worth noting that the result in Theorem 1.4 is not uniform and considers only computable expanding Thurston maps on $\widehat{\mathbb{C}}$. Moreover, Theorem 1.3 can be leveraged to extend the above result to the uniform computability of the equilibrium states of computable expanding Thurston maps on a general topological $2$-sphere $S^{2}$. However, such applications would require a discussion on the computability of visual metrics (see e.g. [BM17, Chapter 8]), which would take us too far astray from the core topic of this article.
The conditions in Theorems 1.1 and 1.3 illustrate some methodological distinctions. Specifically, Theorem 1.1 provides an efficient pathway to proving computability when equilibrium states are unique and entropy functions are regular. In contrast, Theorem 1.3 becomes useful for systems with non-unique equilibrium states or those lacking upper semi-continuous entropy, though it demands more sophisticated analysis. Note that computable functions on some subsets are always continuous on the corresponding subsets. Thus, Theorem 1.3 can be applied to some Borel-measurable dynamical systems (not necessarily continuous on the whole space), for example, Pomeau–Manneville maps with geometric potentials. To discuss the topological pressures and equilibrium states in these Borel-measurable settings, we apply Mike Todd’s definition of them in Subsection 3.2.
In our future investigations, the strategies of Theorem 1.3 will be further extended to study the computability of equilibrium states with time complexity in the following settings:
@@
    Non-uniformly expanding maps on smooth manifolds with Hölder continuous hyperbolic potentials.
@@ 1.2. Strategy and organization
Our approaches reflect two distinct strategies for addressing this challenge. The first approach assumes that the measure-theoretic entropy $h_{\mu}(T)$ coincides with its upper semi-continuous regularization $\overline{h}_{\mu}(T)$ (see Section 4) and uses Lemma 4.6 to characterize $\overline{h}_{\mu}(T)$. The second approach avoids relying on semi-continuity of entropy by instead investigating Jacobians for equilibrium states, a technique motivated by Rokhlin’s formula (cf. Proposition 5.3).
@@
For simplicity, we denote the former (resp. latter) set by $\mathcal{E}_{1}(T,\phi)$ (resp. $\mathcal{E}_{2}(T,\phi)$). Indeed, by some available results in convex analysis (cf. Lemma 4.3 and Propositions 4.5), the set $\mathcal{E}_{1}(T,\phi)$ coincides the set $C(X)^{*}_{\phi,P_{T}}$. Moreover, by our investigations in ergodic theory (cf. Theorem 5.6), the set $\mathcal{E}_{2}(T,\phi)$ can be described by $\mathcal{M}(X,T;Y,J)$ (defined in 1.2), where $Y\subseteq X$ is a Borel set and $J\colon X\rightarrow\mathbb{R}$ is a Borel function satisfying some assumptions (indeed, $\mathcal{E}_{2}(T,\phi)\cap\mathcal{P}(X;Y)$ coincides with $\mathcal{M}(X,T;Y,J)\cap\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$). It is worth noting that the set $\mathcal{E}(T,\phi)$ of equilibrium states may not be weak compact. However, the sets $C(X)^{*}_{\phi,P_{T}}$ and $\mathcal{M}(X,T;Y,J)$ are always weak compact. Hence, we demonstrate the recursive compactness of these sets instead of investigating $\mathcal{E}(T,\phi)$ directly.
@@
Finally, we summarize the two approaches as follows. In the topological pressure approach, we assume the upper semi-continuity of the pressure function to ensure that $\mathcal{E}_{1}(T,\phi)=\mathcal{E}(T,\phi)$. In the Jacobian approach, the main challenge is constructing a recursively compact set $\mathcal{K}\subseteq\mathcal{P}(X)$ to exclude the measures with positive mass on a “bad” region. Since a uniform construction of such a set $\mathcal{K}$ is not feasible for all systems, we hypothesize its existence in Theorem 1.3 to obtain a general result, and provide explicit constructions for specific families in Theorems 5.95.10, and 1.4. More precisely, we consider the case where the “bad” regions are indeed sets of finitely many points. Theorem 5.9 uses a sequence of open sets containing these “bad” regions to eliminate the mass supported on them, Theorem 5.10 investigates the dynamical systems that are uniformly contracting near all the periodic singular points, and Theorem 1.4 studies expanding Thurston maps, a family of maps which are uniformly expanding near all the periodic singular points.
We now describe the structure of the article. After fixing some notation in Section 2, we review some notions and results in computable analysis and ergodic theory in Section 3.
Section 4 focuses on the first approach. We prove Theorem 1.1 by establishing the recursive compactness of the set of tangent functionals to the topological pressure function at the potential with upper semi-continuous entropy.
Section 5 is devoted to the second approach as stated in Theorem 1.3, which uses the prescribed Jacobian with respect to an equilibrium state. Through employing the transfer operators to establish Theorem 5.6 (in Subsection 5.1), which characterizes an equilibrium state in terms of its Jacobian, we complete the proof of Theorem 1.3 in Subsection 5.2. Theorems 5.9 and 5.10, presented as consequences of Theorem 1.3, are stated in Subsection 5.3.
Section 6 examines the computability of the equilibrium states for expanding Thurston maps. We first provide the definitions and properties of these maps in Subsection 6.1. Then in Subsection 6.2, we apply Theorem 1.1 to demonstrate the computability of the equilibrium state for a Misiurewicz–Thurston rational map and a Hölder continuous potential, thereby establishing Theorem 1.2. Finally, Subsection 6.3 addresses a broader class of expanding Thurston maps where entropy map may lack upper semi-continuity. Here, we use Theorem 1.3 to study the measures of maximal entropy and establish Theorem 1.4.
Acknowledgments. Q. He, Z. Li, and X. Shi were partially supported by Beijing Natural Science Foundation (JQ25001 and 1214021) and National Natural Science Foundation of China (12471083, 12101017, 12090010, and 12090015). I. Binder was partially supported by an NSERC Discovery grant. Q. He was also supported by Peking University Funding (7101303303 and 6201001846).

### v2

@@ Abstract.
We investigate the theory of thermodynamic formalism from the perspective of computable analysis, with a special focus on the computability of equilibrium states. Specifically, we develop two complementary general approaches to verify the computability of equilibrium states for nonuniformly expanding computable dynamical systems. The first approach applies to dynamical systems whose topological pressure functions admit effective approximations and whose measure-theoretic entropy functions are upper semicontinuous. As a concrete application, we establish the computability of the equilibrium states for Misiurewicz–Thurston rational maps with Hölder continuous potentials. The second approach exploits prescribed Jacobians of equilibrium states through a local analysis and applies to settings where the measure-theoretic entropy functions may lack upper semicontinuity.
@@ 1. Introduction
Thermodynamic formalism emerges as the natural framework for studying chaotic dynamics from a statistical viewpoint. This theory, which draws inspiration from statistical mechanics, was pioneered by Ruelle, Sinai, Bowen, and others in the 1970s ([Do68, Si72, Bo75, Wa82]). Since its birth, thermodynamic formalism has been extensively applied in various classical contexts and has remained at the frontier and core of research in dynamical systems. It focuses on equilibrium states, which are invariant measures that maximize pressure functionals encoding both entropy [^1] The concepts of entropy in dynamical systems have their roots in the early works on the notions of entropy by Boltzmann and Gibbs (statistical mechanics, 1875), von Neumann (quantum mechanics, 1932), and Shannon (information theory, 1948). These notions of entropy are all designed to describe the complexity of their respective systems or objects. In recent years, there have been exciting developments and diverse applications of entropy and complexity theory, see e.g. Braverman’s report at the International Congress of Mathematicians in 2022 [Bra23]. and integrals of potentials. Crucially, equilibrium states characterize the typical behavior of the dynamics and thus possess central mathematical and physical significance. In many settings, equilibrium states describe the weighted distribution of random backward orbits (see e.g. [HT03, Li18]), iterated preimages, and periodic orbits.
@@
In this article, we study thermodynamic formalism from the point of view of computable analysis, with a special focus on the computability of equilibrium states in dynamical systems. Classical hyperbolic systems admit well-developed techniques and are generally regarded as well-understood. By contrast, nonuniformly hyperbolic systems resist conventional approaches. The relaxation of uniform expansion requirements therefore represents a fundamental challenge in dynamical systems. Our central contribution lies in demonstrating that statistical computability can coexist with dynamical complexity, even in nonuniformly hyperbolic regimes.
@@
The first approach utilizes the set of tangent functionals to the topological pressure function to link the computability of the topological pressure to that of the equilibrium state. This method is suitable for dynamical systems where the measure-theoretic entropy map is upper semicontinuous and the topological pressure function can be effectively approximated. For dynamical systems whose measure-theoretic entropy functions may lack upper semicontinuity, we implement the second approach, which establishes the computability of equilibrium states through local (rather than global) considerations. The novelties of our second approach include the following two aspects: (i) on the thermodynamic formalism side, using general transfer operators to study equilibrium states of full measure away from singular points with prescribed Jacobians, and (ii) on the computable analysis side, the verification of Jacobians away from singular points, and the construction of a recursively compact set of measures to exclude atomic measures supported on singular points.
@@
Based on these works, ergodic theory for expanding Thurston maps has been actively investigated in [BM10, BM17, HP09, Li15, Li17, Li18, LZ25, LSZ25, LS24a, LS24b]. In particular, the third-named author [Li18] developed the thermodynamic formalism and investigated the existence, uniqueness, and other ergodic properties of equilibrium states for expanding Thurston maps333cf. the monograph [Li17].. Notably, expanding Thurston maps exhibit weak expansion; they are neither expansive nor $h$-expansive. Furthermore, those with at least one periodic critical point are not even asymptotically $h$-expansive [Li15]. Recent advances by the third-named and fourth-named authors [LS24b] further demonstrated that the measure-theoretic entropy function is upper semicontinuous if and only if the expanding Thurston map has no periodic critical points. Leveraging these results, we apply our methods to investigate the computability of measures of maximal entropy and equilibrium states for expanding Thurston maps, demonstrating the applicability of our two distinct approaches.
Our approaches to investigating the computability of equilibrium states extend way beyond the setting of expanding Thurston maps. See the discussions below.
@@ 1.1. Main results
The first approach (Theorem 1.1) applies to systems with upper semicontinuous measure-theoretic entropy functions. It shows that certain computability properties of the topological pressures guarantee the computability of the equilibrium states, with a concrete application to Misiurewicz–Thurston rational maps in Theorem 1.2. The second approach, established in Theorem 1.3, applies to systems without relying on the upper semicontinuity of the entropy, where the computability of equilibrium states arises from the computability of their prescribed Jacobians. As an application, we establish the computability of the measures of maximal entropy for computable expanding Thurston maps with computable critical points in Theorem 1.4.
@@
We say that the quintuple $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ is a uniformly computable system if the following properties are satisfied:
@@
In the first approach, we exploit the set of tangent functionals to the topological pressure function $P(T,\cdot)$ (see Section 4) at the potential $\phi$ and demonstrate that the computability of the equilibrium state follows from certain computability properties of the topological pressure. We note that this set of tangent functionals is naturally identified with a subset of the set $\mathcal{P}(X)$ of Borel probability measures (see Remark 4.4). We denote by $\mathcal{E}(T,\phi)$ the set of equilibrium state(s) for a map $T$ and a potential $\phi$. We refer the reader to Subsection 3.1 for a detailed discussion on the computability of measures.
@@
Let $(X,\,\rho,\,\mathcal{S},\,\{X_{n}\}_{n\in\mathbb{N}},\,\{T_{n}\}_{n\in\mathbb{N}})$ be a uniformly computable system with $X_{n}=X$ for all $n\in\mathbb{N}$, and $\{\phi_{n}\}_{n\in\mathbb{N}}$ be a sequence of uniformly computable functions $\phi_{n}\colon X\rightarrow\mathbb{R}$. Suppose $T_{n}$ has finite topological entropy, and the measure-theoretic entropy map $\nu\mapsto h_{\nu}(T_{n})$ is upper semicontinuous on $\mathcal{M}(X,T_{n})$ for each $n\in\mathbb{N}$. Assume that the following statements are true:
@@
Here the space $\mathcal{M}(X,T_{n})$ of $T_{n}$-invariant Borel probability measures is equipped with the weak topology, the computability properties of real numbers are recalled in Definitions 3.3 and 3.6, and the computable structure on the space $\mathcal{P}(X)$ of Borel probability measures is specified in Proposition 3.26.
Applying Theorem 1.1, we establish the computability of the equilibrium states for computable Misiurewicz–Thurston rational maps (i.e., postcritically-finite rational maps without periodic critical points which are computable) and computable Hölder continuous potentials. The existence and uniqueness of the equilibrium state for an expanding Thurston map follow from Theorem 6.3.
@@
[…]
It is worth noting that Theorem 1.1 can be applied to establish the computability of the equilibrium states for a wide range of dynamical systems with a unique equilibrium state and an upper semicontinuous measure-theoretic entropy map, such as rational maps with Hölder continuous hyperbolic potentials (see e.g. [DU91]). Due to space limitations, we focus on the current examples and postpone further investigations to future work.
For many dynamical systems, the measure-theoretic entropy map may not be upper semicontinuous, or verifying this property may be difficult. To overcome this issue, we establish a second approach by considering the prescribed Jacobians for equilibrium states.
@@
By Theorem 5.6, in some settings (cf. Definition 5.2), the set defined in (1.2) describes the set of equilibrium states. Moreover, as applications of Theorem 1.3, in Theorems 5.10 and 5.11, we construct two corresponding families of uniformly recursively compact sets $\{\mathcal{K}_{n}\}_{n\in\mathbb{N}}$ satisfying the additional assumptions for two classes of dynamical systems.
Theorem 1.3 extends the methodologies developed in [BBRY11, Theorem A] and [BHLZ25, Theorem 1.1], introducing new techniques to handle both nonuniform expansion and the possible failure of upper semicontinuity of entropy. Compared to the previous results, our approach significantly broadens applicability by relaxing the computability requirement on prescribed Jacobians to only lower semi-computability.
To demonstrate the applicability of Theorem 1.3 while minimizing unnecessary technicalities, we employ it to prove the computability of the measures of maximal entropy for computable expanding Thurston maps. In fact, one can establish the computability of equilibrium states for these maps with Hölder continuous potentials using the cone method adapted for computability as in [BHLZ25].
@@
We provide examples of such expanding Thurston maps and apply Theorem 1.4 to them at the end of Subsection 6.3. Even though the result in Theorem 1.4 is not uniform and considers only computable expanding Thurston maps on $\widehat{\mathbb{C}}$, Theorem 1.3 can indeed be leveraged to extend the above result to the uniform computability of the equilibrium states of computable expanding Thurston maps on a general topological $2$-sphere $S^{2}$. However, such applications would require a discussion on the computability of visual metrics (see e.g. [BM17, Chapter 8]), which would take us too far astray from the core topic of this article.
The conditions in Theorems 1.1 and 1.3 illustrate some methodological distinctions. Specifically, Theorem 1.1 provides an efficient pathway to proving computability when equilibrium states are unique and measure-theoretic entropy functions are regular. In contrast, Theorem 1.3 becomes useful for systems with nonunique equilibrium states or those (potentially) lacking upper semicontinuous entropy, though it demands more sophisticated analysis. Note that computable functions on some subsets are always continuous on the corresponding subsets. Theorem 1.3 can be applied to some Borel-measurable dynamical systems (not necessarily continuous on the whole space), for example, Pomeau–Manneville maps with geometric potentials.
In fact, Theorem 1.3 can be applied and further extended to study the computability of equilibrium states with time complexity in the following settings:
@@
    Nonuniformly expanding local diffeomorphism on smooth manifolds and Hölder continuous potentials with not very large oscillation (see [BHLS25]).
We leave these investigations for future work.
@@ 1.2. Strategy and organization
Our approaches reflect two distinct strategies for addressing this challenge. The first approach applies to dynamical systems where the measure-theoretic entropy $h_{\mu}(T)$ coincides with its upper semicontinuous regularization $\overline{h}_{\mu}(T)$. Under this hypothesis, we use convex analysis to characterize the set of equilibrium states as the set of tangent functionals to the topological pressure function (see Section 4). The second approach avoids relying on semicontinuity of entropy by instead investigating Jacobians for equilibrium states, a technique motivated by Rokhlin’s formula (cf. Proposition 5.3).
@@
For simplicity, we denote the former (resp. latter) set by $\mathcal{E}_{1}(T,\phi)$ (resp. $\mathcal{E}_{2}(T,\phi)$). Indeed, by results in convex analysis (cf. Lemma 4.3 and Proposition 4.5), the set $\mathcal{E}_{1}(T,\phi)$ coincides with the set $C(X)^{*}_{\phi,P_{T}}$. Moreover, by our investigations in ergodic theory (cf. Theorem 5.6), the set $\mathcal{E}_{2}(T,\phi)$ can be described by $\mathcal{M}(X,T;Y,J)$ (defined in (1.2)), where $Y\subseteq X$ is a Borel set and $J\colon X\rightarrow\mathbb{R}$ is a Borel function satisfying some assumptions (indeed, $\mathcal{E}_{2}(T,\phi)\cap\mathcal{P}(X;Y)$ coincides with $\mathcal{M}(X,T;Y,J)\cap\mathcal{M}(X,T)\cap\mathcal{P}(X;Y)$). It is worth noting that the set $\mathcal{E}(T,\phi)$ of equilibrium states may not be weak compact. However, the sets $C(X)^{*}_{\phi,P_{T}}$ and $\mathcal{M}(X,T;Y,J)$ are always weak compact. Hence, we demonstrate the recursive compactness of these sets instead of investigating $\mathcal{E}(T,\phi)$ directly.
@@
Finally, we summarize the two approaches as follows. In the global approach via convex analysis, we assume the upper semicontinuity of the pressure function to ensure that $\mathcal{E}_{1}(T,\phi)=\mathcal{E}(T,\phi)$. In the local approach via transfer operator and Jacobian, the main challenge is to construct a recursively compact set $\mathcal{K}\subseteq\mathcal{P}(X)$ that excludes measures with positive mass on a “bad” region. Since a uniform construction of such a set $\mathcal{K}$ is not feasible for all systems, we hypothesize its existence in Theorem 1.3 to obtain a general result, and provide explicit constructions for specific families in Theorems 5.105.11, and 1.4. More precisely, we consider the case where the “bad” regions are indeed sets of finitely many points. In Theorem 5.10 we use a sequence of open sets containing these “bad” regions to eliminate the mass supported on them, in Theorem 5.11 we investigate the dynamical systems that are uniformly contracting near all the periodic singular points, and in Theorem 1.4 study expanding Thurston maps, a family of maps which are uniformly expanding near all the periodic singular points.
We now describe the structure of this article. After fixing some notation in Section 2, we review some notions and results in computable analysis and ergodic theory in Section 3.
Section 4 focuses on the first (global) approach. We prove Theorem 1.1 by establishing the recursive compactness of the set of tangent functionals to the topological pressure function at the potential with upper semicontinuous entropy.
Section 5 is devoted to the second (local) approach as stated in Theorem 1.3, which uses the prescribed Jacobian with respect to an equilibrium state. Through employing the transfer operators to establish Theorem 5.6 (in Subsection 5.1), which characterizes an equilibrium state in terms of its Jacobian, we complete the proof of Theorem 1.3 in Subsection 5.2. Theorems 5.10 and 5.11, presented as consequences of Theorem 1.3, are stated in Subsection 5.3.
Section 6 examines the computability of the equilibrium states for expanding Thurston maps. We first provide the definitions and properties of these maps in Subsection 6.1. Then in Subsection 6.2, we apply Theorem 1.1 to demonstrate the computability of the equilibrium state for a Misiurewicz–Thurston rational map and a Hölder continuous potential, thereby establishing Theorem 1.2. Finally, Subsection 6.3 addresses a broader class of expanding Thurston maps whose measure-theoretic entropy maps may lack upper semicontinuity. Here, we use Theorem 1.3 to study the measures of maximal entropy and establish Theorem 1.4.
Acknowledgments. Q. H., Z. L., and X. S. were partially supported by Beijing Natural Science Foundation (JQ25001 and 1214021) and National Natural Science Foundation of China (12471083, 12101017, 12090010, and 12090015). I. B. was partially supported by an NSERC Discovery grant. Q. H. was also supported by Peking University Funding (7101303303 and 6201001846).